United Snack Company sells 50-pound bags of peanuts to university dormitories for $20 a bag. The fixed costs of this operation are $176,250, while the variable costs of peanuts are $0.15 per pound a. What is the break-even point in bags? b. Calculate the profit or loss on 7,000 bags and on 20,000 bags. c. For next year the company assumes all costs and the sales price to remain the same, but wants to generate additional gross profit of $50,000. What would be the new break-even point in bags for next year?
Answer #1

Ans A)

Sales Price = 20 Per Bag

Variable Price = 0.15 per pound

Pound required per bag = 50

Variable cost per Bag = 50 * 0.15 = 7.5

Break Even Point = Fixed Cost / Contribution Per Unit

where Contribution = Sales Price - Variable Cost

So Break Even Point = 176250/ 20-7.5

= 176250/12

= 14100 Bags

Ans B)

Particulars 7000 Bags 20000 Bags
a Sales Unit 7000 20000
b Sales Price 20 20
c Sales Revenue (a * b) 140000 400000
d Variable Cost PU 7.5 7.5
e Total Variable Cost ( a* d) 52500 150000
f Contribution ( c - e) 87500 250000
g Fixed Cost 176500 176500
h Profit/ -Loss (f - g) -89000

73500

Ans C)

For Next Year When they want additional GP of 50000 we need to increase the Profit Margin by temporary increasing Fixed Cost so as to calculate new Break Even

New Fixed Cost = 176250 + 50000 = 226250

Break Even Point = Fixed Cost/ Contribution PU

= 226250/12.5

= 18100 Units